# Viewpoint Matters: Dynamically Optimizing Viewpoints with Masked Autoencoder for Visual Manipulation

### v1

@@ 2.3 Next Better Viewpoint Selection
Fig. 2: Visualization of the selected viewpoints in our experiments, showcasing both simulation and real-world environments. Each row represents the procedure of a specific task, indicating the necessity of selecting different viewpoints.
3 Experiments
We evaluate MAE-Select across 3 challenging scenarios and 8 demanding tasks in simulation, including simulations in ACT [25], RLBench [11], and our designed robot in MuJoCo [22], along with three real-world tasks.
3.1 Implementation
@@
We explore a single-camera control setup where the system is trained with multiple camera views but deployed with a single moving camera. At each time chunk, the camera is placed at one of the training viewpoints, enabling flexible operation in real-world settings where only one camera is available. We base our implementation on the architecture of the diffusion policy [3]. The action space corresponds to the joint angles of the robot arm, while the image observations have a resolution of $224\times 224$ with a patch size of 16. Our masked autoencoder utilizes a 12-layer ViT [4] for encoder and an 8-layer ViT for decoder, with an embedding dimension of 512. During pretraining, we use a batch size of 128 over 100 epochs. For RLBench and real-world tasks, the time chunk size is set to 20, with a total of 600 epochs. For other tasks, the time chunk size is 100, with 1,000 epochs in total. $\lambda_{1}$ and $\lambda_{2}$ are empirically set to 2.0 and 1.0, respectively. For each method, we evaluate the best-performing checkpoints from the last three evaluated at 100-intervals, with 50 environment initializations in simulation. All models were trained and tested on NVIDIA RTX 4090 GPUs.
3.2 Results
We compare MAE-Select with Diffusion Policy [3] and its MAE-enhanced variant (MAE-Diffusion) across various tasks. Unlike the baselines, which use fixed viewpoints during training and testing, MAE-Select leverages multi-view training and dynamically selects the most informative single viewpoint for the next time chunk during inference.
As shown in Tab.1, MAE-Select consistently outperforms fixed single-camera setups in both simulation and real-world experiments—for instance, achieving an 8% gain over the best fixed camera method and 32% over prior work in the Put Box In Cabinet task. Its strength lies in intelligently exploiting limited visual input through optimal viewpoint selection. Interestingly, in some tasks a single camera even surpasses multi-camera setups (e.g., Diffusion Policy on Unplug Charger: top view 44% vs. two views 34%), likely due to noise and alignment issues from multiple inputs. MAE-Select avoids such pitfalls, remaining competitive and often superior, even compared with multi-camera configurations.
To analyze system behavior under single-camera control, we visualize deployment phases (Fig.2). MAE-Select adaptively selects optimal viewpoints, focusing on critical regions while ignoring irrelevant areas.
3.3 Ablation Study
Our viewpoint selection method is agnostic to the choice of action decoder, as evidenced by its integration with ACT [25] (Tab.2), confirming its adaptability across architectures. We further conduct an ablation study comparing the full encoder-decoder masked autoencoder with a variant using only the encoder [20, 21]. Results (Tab.3) show that leveraging the full structure significantly boosts performance, especially under occlusion, highlighting the decoder’s role in refining visual representations and improving generalization in both simulation and real-world tasks.
@@ 4 Conclusions and Future Work
We propose MAE-Select, a framework that enables single-camera robots to actively optimize viewpoints for manipulation tasks. By exploiting pre-trained multi-view masked autoencoder representations and dynamically selecting informative viewpoints at each time chunk without manual annotations, MAE-Select boosts manipulation efficiency and can even outperform multi-camera systems. Its main limitation is optimizing over discrete rather than continuous viewpoints, which limits flexibility in dynamic environments. Future work could integrate NeRF [15] or 3D Gaussian processes [13] for continuous viewpoint optimization.

### v2

@@ 4.3 Implementation
4.4 Results
We compare the performance of MAE-Select with two other methods: Diffusion Policy [chi2023diffusion] and a variant of Diffusion Policy that incorporates MAE, referred to as MAE-Diffusion, across different types of tasks and viewpoints. For Diffusion Policy and MAE-Diffusion, the observation viewpoint configuration is held consistent between training and testing phases. In contrast, MAE-Select employs a multi-view training strategy that incorporates both available viewpoints, while adopting a dynamic viewpoint selection mechanism during testing where only a single optimal viewpoint is activated per time chunk.
As demonstrated in Tab.1, MAE-Select consistently outperforms other fixed single-camera setups in both simulation and real-world experiments. For example, in the Put Box In Cabinet task, MAE-Select improves performance by 8% compared to the best fixed single-camera method and by 32% compared to previous work. Its advantage lies in its ability to intelligently select the most informative viewpoints, which allows the system to make the most of limited visual input, resulting in optimized task completion.
Furthermore, an interesting pattern emerges in some tasks: for certain methods, the performance with a single viewpoint can surpass that of a multi-camera setup. For example, in the Unplug Charger task with Diffusion Policy, using only the top view (44%) outperforms using both views (34%). This counterintuitive result may stem from the added complexity of processing multiple cameras, which can introduce noise or misalignment issues, complicating the learning process. By focusing on the optimal viewpoint, MAE-Select avoids these challenges, enabling more efficient and effective task execution. Consequently, MAE-Select remains highly competitive when compared to multi-camera setups, even outperforming them in several tasks.
4.5 Ablation Study
We conduct a comprehensive series of ablation studies to validate the efficacy and necessity of the core components within our proposed MAE-Select architecture. This section primarily investigates the framework along two dimensions: compatibility with various action decoders and the utilization efficiency of the Masked Autoencoder (MAE) structure.
@@ 4.5 Ablation Study
4.5.1 Action Decoder Compatibility
A primary advantage of our method is the high decoupling of its viewpoint selection mechanism from the specific downstream action decoder. This implies our viewpoint selection module can serve as a perception component, flexibly adapting to various action generation models. To highlight this versatility, we also evaluate our approach in combination with ACT [zhao2023learning], an alternative action decoder. As shown in Table 2, integrating our method significantly enhances baseline architectures. Notably, in the Phone On Base task, MAE-Select reaches a 70% success rate, outperforming both ACT (58%) and MAE-ACT (66%). This demonstrates that dynamic single-view selection can integrate seamlessly into various architectures, providing a sharper decision basis and lower input redundancy than fixed multi-view fusion.
@@ 4.5.1 Action Decoder Compatibility
4.5.2 Masked Autoencoder
We also conduct an ablation study to isolate the contributions of fully utilizing the entire encoder-decoder structure of the masked autoencoder in our approach. Specifically, we compare the performance of our method, which leverages both the encoder and decoder components, against a version that uses only the encoder of the masked autoencoder [seo2023masked, seo2023multi]. The results in Tab.3 demonstrate that utilizing the full encoder-decoder structure significantly improves performance, particularly in scenarios that require nuanced visual understanding from partial or occluded viewpoints. This highlights the value of the decoder in refining the system’s ability to interpret and act based on incomplete or masked information, contributing to better generalization and adaptability in both simulated and real-world tasks.
Figure 5: Visualization of the selected viewpoints in our simulation environments. The unmasked images represent the specific views selected by the model.
Figure 6: Visualization of the selected viewpoints in our real-world environments. Each row represents the procedure of a specific task, indicating the necessity of selecting different viewpoints.
4.6 Visualization
To intuitively understand the decision-making logic and behavioral patterns of the system under the single-camera control setup, we provide a comprehensive set of visualizations extracted during the deployment phases. As illustrated in Fig. 5 and Fig. 6, MAE-Select exhibits highly intelligent, human-attention-like decision-making capabilities—it dynamically and accurately selects the optimal viewpoint based on the task stage and contextual information. In the simulation environments (see Fig. 5), we observe the temporal progression of the Unplug Charger and Put Box In Bin with Disturbance tasks. Taking Unplug Charger as an example: in the early stages of the task (T=0 to T=20), the system selects the third-person viewpoint, a global perspective to establish the relative spatial arrangement of the robotic arm, charger, and socket. However, as the end-effector gradually approaches the target and requires high-precision alignment and extraction (T=40 and beyond), the system intelligently selects the wrist viewpoint. This dynamic transition effectively mitigates severe occlusions caused by the robot’s own body, highlighting MAE-Select’s ability to prioritize critical interaction areas while disregarding irrelevant background regions.
@@ 5 Conclusions and Future Work
In this work, we present MAE-Select, a novel framework that enables a robot to actively optimize its viewpoint for manipulation tasks using only a single camera. By fully leveraging pre-trained representations from multi-view masked autoencoders and dynamically selecting the next most informative viewpoints at each time chunk without manual annotations, MAE-Select significantly enhances the efficiency of robotic manipulation, addressing the limitations of both multi-camera and single-view setups. Our experiments demonstrate that this novel mechanism effectively improves performance, and even surpasses multi-camera systems in certain cases. Despite its effectiveness, one major limitation of MAE-Select is that it optimizes over discrete viewpoints rather than continuous ones, which reduces the system’s flexibility in dynamic environments. Future improvements could involve the integration of techniques like Neural Radiance Fields (NeRF) [mildenhall2021nerf] or 3D Gaussian processes [kerbl20233d], enabling continuous viewpoint optimization.